Washington State Fisheries, Inc., processes salmon for various distributors. Two departments, processing and packaging, are involved. Data relating to tons of salmon sent to the processing department during May follow:

Percent Completed
Tons of Salmon Direct Materials Conversion
Work-in-process inventory, May 1 1,900 80 % 70 %
Work-in-Process inventory, May 31 2,040 50 % 30 %
Started processing during May 10,300

Required:

1. Calculate the number of tons completed and transferred out during the month.

2. Calculate the number of equivalent units for both direct materials and conversion for the month of May, assuming that the company uses the weighted-average method.

3. How would your answer in requirement 2 change if the percentage of completion in ending inventory were as follows: direct materials 30%, conversion 40%?
Answer #1

1. Units completed and transferred out

= Opening WIP + started processing - closing WIP

= 1900 + 10300 - 2040

= 10160

2. In weighted average method , opening units will be considered as completed in current year only hence 100% will be counted towards them.

Calculation of equivalent units

Particular Material conversion cost
% units % units
Units transferred out 100 10160 100 10160
Finished goods WIP 50% 1020 30% 306
Equivalent units 11180 10466

3. Only finished goods will be changed.

New material in finished goods = 1020*30% = 306

Conversion cost = 1020*40% = 408

So new equivalent units will be

Material = 10160 + 306 = 10466

Conversion cost = 10160 + 408 = 10568

Feel free to ask any queries..

Also plz upvote it means a lot .. thank you